# From Flat to Hierarchical : Extracting Sparse Representations with Matching Pursuit

### v1

@@ B.4 Multimodal Models
To quantify the extent to which learned features specialize by modality or capture shared structure, we compute the Modality Score for each feature, following the formulation of [105]. Values near $1$ indicate image-specific features; values near 00 indicate text-specific features; and intermediate values reflect balanced, multimodal activation. To ensure that Modality Scores reflect relative activation patterns rather than absolute energy differences between modalities, we scale the energy of text inputs by a factor of $1/5$ before computing the above quantities. This corrects as we have 5 times more captions than images. We find that the same overall patterns emerge when we apply a modality wise normalization.
@@
B.5 Preliminary Experiments with Language Models using TinyStories
@@ B.5.1 Experimental details
To assess whether our analysis of the vision domain (see Sec. B.4) generalizes to a language modeling setting, we perform preliminary experiments using a 4-layer Transformer model [144] pretrained on the TinyStories dataset [145]. We train the Vanilla [14], TopK [15], and MP-SAE (see Sec. 3) on residual stream representations extracted from midway through the model, i.e., end of the second Transformer block (from amongst 4 total blocks). $p$, which refers to the width of our SAEs, i.e., the size of the sparse code $z$, is set to be $4\times m$ for all SAEs; here, $m$ corresponds to the dimensionality of the residual stream. For TopK and MP-SAE, we set $\ell_{0}=100$ for any inputted representation; for Vanilla SAE, we search for a value of $\lambda$ to ensure after training $\ell_{0}$, on average, is $100$. Inline with prior work [22], activations are standardized by computing a population mean and standard deviation and rescaled to be, on expectation, $\sqrt{m}$ magnitude in norm. Training is performed using Adam optimizer [146], with a constant learning rate of $10^{-3}$, $\beta_{1},\beta_{2}=0.9,0.95$, weight decay of $10^{-4}$, and gradient clipping at unit-norm magnitude for all the SAEs. Batch-size is set to be $5000$ tokens, each of which is randomly sampled from samples (stories) from the train-split of TinyStories. Training goes on for approximately $40$K iterations. All results in the following experiments are performed on samples drawn from the eval-split of TinyStories.
@@ B.5.2 Results
Similar to Fig. 8, for all SAEs, we take their parameters trained up to iteration $t$ and compare the inputted representations ($x$) with the ones predicted by the SAE given its $k$ largest magnitude latents in the sparse code, denoted $x_{t,k}$. Specifically, we compute the normalized MSE, $\mathbb{E}_{x}\left[\nicefrac{{||\hat{x}_{t,k}-x||^{2}}}{{||x||^{2}}}\right]$, for increasing values of $t$ and $k$. We note this experiment is different from the one reported in Fig. 8 because it also assesses the effects of training, hence going beyond merely highlighting the inference-time flexibility of MP-SAE. Such an analysis is manageable in the current setting since models are not exceptionally big ($\sim 1$M parameters).
@@ Appendix C Theoretical Guarantees for MP-SAE Inference
Let ${\bm{r}}^{(t)}$ denote the residual at iteration $t$ of MP-SAE inference, and let $j^{(t-1)}$ be the index of the feature selected at step $t{-}1$. If the column $j^{(t-1)}$ of the dictionary $\bm{D}$ satisfy $\|\bm{D}_{j^{(t-1)}}\|_{2}=1$, then the residual becomes orthogonal to the previously selected feature:

### v2

@@ B.4 Multimodal Models
To quantify the extent to which learned features specialize by modality or capture shared structure, we compute the Modality Score for each feature, following the formulation of [105]. Values near $1$ indicate image-specific features; values near 0 indicate text-specific features; and intermediate values reflect balanced, multimodal activation. To ensure that Modality Scores reflect relative activation patterns rather than absolute energy differences between modalities, we scale the energy of text inputs by a factor of $1/5$ before computing the above quantities. This corrects as we have 5 times more captions than images. We find that the same overall patterns emerge when we apply a modality wise normalization.
@@
B.5 Preliminary Experiments with Language Models
@@ B.5.1 Experimental details
To assess whether our analysis of the vision domain (see Sec. B.4) generalizes to a language modeling setting, we perform preliminary experiments using a 4-layer Transformer model [145] pretrained on the TinyStories dataset [146]. [^6] We also provide a trained MP-SAE on Gemma activations at https://github.com/eslubana/mpsae. We train the Vanilla [14], TopK [15], and MP-SAE (see Sec. 3) on residual stream representations extracted from midway through the model, i.e., end of the second Transformer block (from amongst 4 total blocks). $p$, which refers to the width of our SAEs, i.e., the size of the sparse code $z$, is set to be $4\times m$ for all SAEs; here, $m$ corresponds to the dimensionality of the residual stream. For TopK and MP-SAE, we set $\ell_{0}=100$ for any inputted representation; for Vanilla SAE, we search for a value of $\lambda$ to ensure after training $\ell_{0}$, on average, is $100$. Inline with prior work [22], activations are standardized by computing a population mean and standard deviation and rescaled to be, on expectation, $\sqrt{m}$ magnitude in norm. Training is performed using Adam optimizer [147], with a constant learning rate of $10^{-3}$, $\beta_{1},\beta_{2}=0.9,0.95$, weight decay of $10^{-4}$, and gradient clipping at unit-norm magnitude for all the SAEs. Batch-size is set to be $5000$ tokens, each of which is randomly sampled from samples (stories) from the train-split of TinyStories. Training goes on for approximately 4040K iterations. All results in the following experiments are performed on samples drawn from the eval-split of TinyStories.
@@ B.5.2 Results
[…]
@@ Appendix C Theoretical Guarantees for MP-SAE Inference
Let ${\bm{r}}^{(t)}$ denote the residual at iteration $t$ of MP-SAE inference, and let $j^{(t-1)}$ be the index of the feature selected at step t1t{-}1. If the column $j^{(t-1)}$ of the dictionary $\bm{D}$ satisfy $\|\bm{D}_{j^{(t-1)}}\|_{2}=1$, then the residual becomes orthogonal to the previously selected feature: